OpenStudy (anonymous):

What is the ratio of the perimeter of the larger rectangle to the smaller rectangle if the perimeter of the larger rectangle is 24 and the perimeter of the smaller rectangle is 21. The options are 11:9, 3:4, 4:3, and 5:3.

OpenStudy (anonymous):

Let P be the perimeter of the larger rectangle and p be the perimeter of the smaller rectangle. Then the ratio P:p = 24:21 = 8:7 Am I missing something?

OpenStudy (anonymous):

Here is the original question (I thought I had the perimeters figured out, but I could have been wrong): the dimensions of a smaller rectangle are 3 ft. by 9 ft. The dimensions of a larger rectangle are 5 ft. by 11 ft. Find the ratio of the perimeter of the larger rectangle to the perimeter of the smaller rectangle. Options are 11:9, 3:4, 4:3, and 5:3

OpenStudy (anonymous):

Then p= 2*(3+9) = 24 ft. and P = 2*(5+11) = 32 Now, P:p = 32:24 = 4:3